OpenStudy (anonymous):

What is one root of this equation? 2x2 - 4x + 9 = 0

OpenStudy (danjs):

have you done factoring in your class, making that equation look like (x )(2x ) = 0 ?

OpenStudy (danjs):

or the use of the quadratic formula with a = 1 b = -4 c = 9

OpenStudy (mathmate):

@onemoretest Have you covered complex numbers?

OpenStudy (anonymous):

i have done it, i just font remember the process. so i would plug one of those three numbers into this? (x )(2x ) = 0 and i don't remember doing that @mathmate

OpenStudy (danjs):

A failsafe method is just to use the quadratic formula the \[x = \frac{ -b +- \sqrt{b^2 - 4*a*c} }{ 2a }\]

OpenStudy (mathmate):

If you have already covered complex numbers, then use the quadratic formula for \(2x^2-4x+9\) \(\huge x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) where a=2, b=-4 and c=9. You will end up with an expression involving i.

OpenStudy (anonymous):

thank you :) i got it

OpenStudy (mathmate):

If you were referring to me or both, you're welcome!